QB=1691-400PB+6pA-6y
PB=0.1 pA=0.3 y (INCOME)=10
1.PRICE ELASTICITY OF DEMAND
2.INCONE ELASTICITY OF DEMAND
1
Expert's answer
2021-08-09T11:28:50-0400

 "QB = 1691 - 400\u00d70.1 + 6\u00d70.3 - 6\u00d710 = 1646.8" units.

A) Price elasticity of demand for the good is: "Ed =\frac{ -400×0.1}{1646.8} = -0.024," so the demand is inelastic.

B) Income elasticity of demand for good B is:

"Ei =\frac{ -6×10}{1646.2} = -0.036," so the good is inferior.